OpenStudy (anonymous):

I wish to double check this answer! A motor boat travels in a straight path on a calm lake with a speed of 7.4 m/s. suddenly the motor stops and the boat starts to slow at a rate of .9 m/s^2. Determine the distance it takes for the boat to come to a stop.

OpenStudy (anonymous):

I went with 0=.9(t) + 7.4 to solve for time = 8.2 then stuck it in x=1/2(.9)(8.2)^2 + 7.4(8.2) + 0 (?) = 96.94m does anyone know if this looks right?

OpenStudy (abmon98):

so u=7.4 v=0 a=.9 so apply the equation that says= v^2-u^2/2a

OpenStudy (abmon98):

-(7.4)^2/2(.9) will give you the distance

OpenStudy (anonymous):

ah thanks! I just made .9 negative because it was slowing

OpenStudy (abmon98):

your most welcome :)